OpenStudy (anonymous):

Write the equation of a line that is parallel to the line 2x - 3y = 5 and passes through the point (2,-1).

OpenStudy (anonymous):

Gradient will be the same, so put the constant as c, plug in the new coordinates and you'll get c, reform the equation, give it a shot :) (Use y = mx + c and make y the subject of the equation)

OpenStudy (anonymous):

is it 3x+2y=5, 3x-2y=7,2x-3y=7, or 2x+3y=1?

OpenStudy (texaschic101):

you first need to find the slope of the equation given. Do you know how to do that ?

OpenStudy (anonymous):

i got 3y -2x = -7, as y = 2/3x - 5/3, take 5/3 = c, you get y = 2/3 x + c, and then replace x and y with 2, -1. Solve it you get c = -7/3 now plug it all in. :P

OpenStudy (texaschic101):

2x - 3y = 5 -3y = -2x + 5 y = 2/3x - 5/3 slope is 2/3. A parallel line will have the same slope. now we will use y = mx + b slope(m) = 2/3 (2,-1) x = 2 and y = -1 now sub into y = mx + b and solve for b -1 = 2/3(2) + b -1 = 4/3 + b -1 - 4/3 = b -3/3 - 4/3 = b -7/3 = b so your parallel equation is : y = 2/3x - 7/3 but you need it in standard form. Ax + By = C y = 2/3x - 7/3 --multiply by 3 3y = 2x - 7 -- subtract 2x -2x + 3y = -7 -- multiply ny -1 to make x positive 2x - 3y = 7 mathbreaker is correct :)